# DAMS-DETR: Dynamic Adaptive Multispectral Detection Transformer with Competitive Query Selection and Adaptive Feature Fusion

### 01 Mar 2024 v1

@@ 3.1 Overview
The overview of our DAMS-DETR is illustrated in Fig. 2. Our method contains four main components: two modality-specific CNN backbones, two modality-specific Efficient Hybrid Encoders,
@@
Modality Competitive Query Selection, and a Multispectral Transformer Decoder. Given a pair of infrared and visible images, we first extract and encode their features separately using two modality-specific CNN backbones and two modality-specific Efficient Hybrid Encoders. Subsequently, the encoded features are flattened, concatenated, and input into the Modality Competitive Query Selection module. This module selects salient modality features to serve as the initial object queries. Next, these modality-specific object queries enter the Multispectral Transformer decoder, which refines them with multiple semantic levels of infrared and visible feature maps through cascaded decoder layers. Finally, these refined object queries are mapped through the detection head to obtain the bounding boxes and classification scores of all objects.
@@ 3.2 Modality Competitive Query Selection
Concretely, we concatenate encoded feature sequences from the infrared and visible modalities and feed them into a linear projection layer to obtain feature point scores. From this combined feature representation,
@@
we select the Top-$K$ scoring features as the initial object queries. These Top-$K$ features are sourced either from the infrared or visible features, respectively, and each feature represents an object instance specific to their respective modality. This approach can be defined as follows:
@@ 3.3 Multispectral Transformer Decoder
Figure 5: Visualization of feature sampling at different semantic levels in different decoder layers. We use different colors of points to represent the results of sampling in different semantic layers, where blue, green, and red represent sampling points on low-level, middle-level, and high-level semantic features maps respectively. Brightly colored and large points indicate relatively high attention weights. (a) The green boxes in the visible image represent aligned bounding boxes, which show the sampling points in each modality are concentrated on the right instance locations. (b) Objects occluded by background and smoke tend to predominantly focus on the infrared modality at each decoder layer, whereas undisturbed objects simultaneously attend to complementary information from both modalities.
@@ 4.3 Comparison with SOTA
Effect of Multispectral deformable cross-attention module (MDCA). We apply standard deformable cross-attention to the fused features obtained by adding the outputs of two encoders in the decoder.
@@
As shown in the 5th and 7th rows of the Tab. 5, our method brings improvements of 1.5% $m$AP50 and 0.9% $m$AP. This improvement can be attributed to our method’s ability of effectively adapting to modality misalignment scenes while performing adaptive feature aggregation for each object in a more fine-grained way. Certainly, when using both Modality Competitive Query Selection and the Multispectral Deformable Cross-Attention module, we achieved the best results on all AP metrics, as shown in the 9th row of the Tab. 5.

### 04 Mar 2024 v2

@@ 3.1 Overview
The overview of our DAMS-DETR is illustrated in Fig. 2. Our method contains four main components: two modality-specific CNN backbones, two modality-specific Efficient Hybrid Encoders, Modality Competitive Query Selection, and a Multispectral Transformer Decoder.
@@
Given a pair of infrared and visible images, we first extract and encode their features separately using two modality-specific CNN backbones and two modality-specific Efficient Hybrid Encoders. Subsequently, the encoded features are flattened, concatenated, and input into the Modality Competitive Query Selection module. This module selects salient modality features to serve as the initial object queries. Next, these modality-specific object queries enter the Multispectral Transformer decoder, which refines them with multiple semantic levels of infrared and visible feature maps through cascaded decoder layers. Finally, these refined object queries are mapped through the detection head to obtain the bounding boxes and classification scores of all objects.
@@ 3.2 Modality Competitive Query Selection
Concretely, we concatenate encoded feature sequences from the infrared and visible modalities and feed them into a linear projection layer to obtain feature point scores. From this combined feature representation, we select the Top-$K$ scoring features as the initial object queries.
@@
These Top-$K$ features are sourced either from the infrared or visible features, respectively, and each feature represents an object instance specific to their respective modality. This approach can be defined as follows:
@@ 3.3 Multispectral Transformer Decoder
Figure 5: Visualization of feature sampling at different semantic levels in different decoder layers. We use different colors of points to represent the results of sampling in different semantic layers, where blue, green, and red represent sampling points on low-level, middle-level, and high-level semantic features maps respectively. Brightly colored and large points indicate relatively high attention weights. (a) The green boxes in the visible image represent aligned bounding boxes, which show the sampling points in each modality are concentrated on the right instance locations. (b) Objects occluded by background and smoke tend to predominantly focus on the infrared modality at subsequent decoder layers. (c) Objects in good illumination conditions and those less distinguishable in the infrared modality tend to predominantly focus on the visible modality at subsequent decoder layers.
@@ 4.3 Comparison with SOTA
Comparisons on VEDAI: As shown in Tab. 4, our method still can work well for small objects, and outperforms most of the SOTA methods. Especially, on $m$AP50, our method outperforms the best (LRAF-Net [5]) by 5.6%. On $m$AP, our method does not outperform LRAF-Net [5] which is a CNN-based detector. The reason could be that in comparison to long-range information extraction, small object detection more relies on local feature extraction, and CNNs could be more advantageous than Transformer in this regard. On the other hand, the smaller bounding boxes of these small objects are sensitive to displacements, and even a few pixel displacements will lead to significant changes in IOU and subsequent $m$AP reduction. Therefore, high precision in bounding box prediction is crucial for small object detection. Relatively, CNN methods can achieve higher bounding box accuracy, resulting in a higher $m$AP.
@@
Nevertheless, the improvement in $m$AP50 demonstrates the potential of our method in small object scenes.
@@
Effect of Multispectral deformable cross-attention module (MDCA). We apply standard deformable cross-attention to the fused features obtained by adding the outputs of two encoders in the decoder. As shown in the 5th and 7th rows of the Tab. 5, our method brings improvements of 1.5% $m$AP50 and 0.9% $m$AP.
@@
This improvement can be attributed to our method’s ability of effectively adapting to modality misalignment scenes while performing adaptive feature aggregation for each object in a more fine-grained way. Certainly, when using both Modality Competitive Query Selection and the Multispectral Deformable Cross-Attention module, we achieved the best results on all AP metrics, as shown in the 9th row of the Tab. 5.